OpenStudy (anonymous):

Hi every buddy, could you please complete this fill in the blank, "Slope intercept form of 5x-12y+39=0 is ___________"

OpenStudy (cwrw238):

12y = 5x + 39 y = (5/12)x + 39/12

OpenStudy (anonymous):

Thanks buddy but shouldn't their some kind of a title for this in the blank or thats the right answer.I m weak in Maths so plz don't mind.

OpenStudy (cwrw238):

no need for a title the general form is y = mx + c where m = slope and c = intercept om y-axis

OpenStudy (cwrw238):

here the slope = 5/12 and intercept is 39/12

OpenStudy (anonymous):

so the actual answer I shoud write in exam is y = (5/12)x + 39/12